OpenStudy (braydonlevi99):

A student has some $1 bills and $5 bills in his wallet. He has a total of 15 bills that are worth $47. How many of each type of bill does he have?

OpenStudy (cp123):

7 of them are 1s and 8 of them are 5s.

OpenStudy (cp123):

47-7 = 40 40/5= 8. so you have 7 ones and 8 fives

OpenStudy (braydonlevi99):

I need better work...

OpenStudy (cp123):

well I just estimated, you should say you did that. I dont know what the actual solution would be

OpenStudy (braydonlevi99):

I'm not writing that down.. Sorry.

OpenStudy (shaima):

so its 8 five dollar bills and 9 one dollar bills??

OpenStudy (braydonlevi99):

No, it's 7 $1 bills and 8 $5 bills. But, I don't know how to show my work.

OpenStudy (shaima):

wait no There are 7 $1 bills and 8 $5 bills.

OpenStudy (shaima):

oh

OpenStudy (cp123):

did you learn in school to estimate or did you learn an actual solution?

OpenStudy (braydonlevi99):

The actual solution.

OpenStudy (shaima):

so n = 15 - m = 15 - 8 = 7

OpenStudy (cp123):

oh ok, sorry I dont know then. what grade is this?

OpenStudy (braydonlevi99):

11th Grade - Algebra II

OpenStudy (cp123):

oh ok i am only in algebra 1

OpenStudy (will.h):

x + y = 15 x + 5y = 47 use substitution method x = -y + 15 x + 5y = 47 now substitue x in the 2nd equation (-y + 15) + 5y = 47 4y + 15 = 47 4y = 32 y = 8 now let's find x x + y = 15 we know that y = 8 x + 8 = 15 x = 7 Now we know we have 7 1s and 15 5s make sense?

OpenStudy (rjtrain1):

top one

OpenStudy (will.h):

lol Thx

OpenStudy (cp123):

u mean 8 5s, since y = 8

OpenStudy (will.h):

oh yes 8 5s and and 7 1s

OpenStudy (braydonlevi99):

Thank you, @Will.H !

OpenStudy (will.h):

sure

OpenStudy (will.h):

let me know if you have other questions i have recently completed alg 2

OpenStudy (braydonlevi99):

Will do!